OpenStudy (anonymous):

If f(x)=2x2−4x−3 and g(x)=2x2−16, find f(x)+g(x).

OpenStudy (anonymous):

put a plus sign between them then combine like terms

OpenStudy (anonymous):

between f(x) and g(x)?

OpenStudy (anonymous):

Solve (2x2−4x−3) + (2x2−16)

OpenStudy (anonymous):

4x-19, is that the answer?

OpenStudy (anonymous):

the problem is 2x^2-4x-3 and 2x^2-16

OpenStudy (welshfella):

4x - 19 is not the answer yiu should add the 2x^2 ' s as well and its - 4x not 4x

OpenStudy (anonymous):

-4x-19?

OpenStudy (anonymous):

or adding the 2x^2's, would it be 4x^2-4x-19?

OpenStudy (texaschic101):

yes.....it would be 4x^2 - 4x - 19

OpenStudy (anonymous):

If f(x)=(x+1)^−1 and g(x)=x−2, what is the domain of f(x)÷g(x)? @texaschic101